# CIF-Bench: A Chinese Instruction-Following Benchmark for Evaluating the Generalizability of Large Language Models

### v1

@@ Abstract
The advancement of large language models (LLMs) has enhanced the ability to generalize across a wide range of unseen natural language processing (NLP) tasks through instruction-following. Yet, their effectiveness often diminishes in low-resource languages like Chinese, exacerbated by biased evaluations from data leakage, casting doubt on their true generalizability to new linguistic territories. In response, we introduce the Chinese Instruction-Following Benchmark (CIF-Bench), designed to evaluate the zero-shot generalizability of LLMs to the Chinese language. CIF-Bench comprises 150 tasks and 15,000 input-output pairs, developed by native speakers to test complex reasoning and Chinese cultural nuances across 20 categories. To mitigate evaluation bias, we release only half of the dataset publicly, with the remainder kept private, and introduce diversified instructions to minimize score variance, totaling 45,000 data instances. Our evaluation of $28$ selected LLMs reveals a noticeable performance gap, with the best model scoring only 52.9%, highlighting the limitations of LLMs in less familiar language and task contexts. This work aims to uncover the current limitations of LLMs in handling Chinese tasks, pushing towards the development of more culturally informed and linguistically diverse models with the released data and benchmark [^1] https://yizhilll.github.io/CIF-Bench/.
@@ 1 Introduction
The landscape of natural language processing (NLP) has been dramatically reshaped by the emergence of large language models (LLMs), which have demonstrated an ability to generalize across unseen NLP tasks, often showcased through the framework of instruction-following Mishra et al. (2021); Sanh et al. (2021); Wei et al. (2021). Despite these advances, skepticism remains regarding the transferability of this instruction-following capability, particularly in multilingual contexts. The models perform worse when switching to Chinese due to the prevalence of English training data Huang et al. (2023b); Zhang et al. (2023b), as figured in Fig. 1. This concern is exacerbated by observations that benchmarks designed to assess the capabilities of LLMs may inadvertently suffer from biased evaluations due to data leakage Sainz et al. (2023), particularly when web-scale datasets are employed to enhance model generalizability Raffel et al. (2023). Such observations raise a critical question: While the generalizability of LLMs appears intriguing, do these models face significant challenges when evaluated in less-common language contexts with newly curated instruction-formatted tasks?
@@ 2 Related Work
Large-scale pre-trained language models have been found that they can generalize across unseen tasks by fine-tuned on formatted task instructions (Khashabi et al., 2020; Mishra et al., 2021; Wei et al., 2021; Sanh et al., 2021). Early studies attempt to fine-tune and evaluate such a capability in a few-shot manner by providing input-output examples Ye et al. (2021); Mishra et al. (2021). Following that, another line of research Bach et al. (2022); Wang et al. (2022b) improves the evaluation reliability from the perspective of scaling the task quantity and providing well-defined corresponding instructions. A more recent concurrent work FollowBench Jiang et al. (2023) proposes to use a similar that crafting multiple instructions for a task to evaluate the LLMs. A core distinction between CIF-Bench and this strategy is that whilst the FollowBench aims at applying different kinds of constraints in the same task, we focus on assessing whether models can stably perform given diversely expressed instructions with the same semantics.
@@ 3 The Challenging Chinese Instruction-Following Benchmark
We define a total of 150 curated tasks, constructed according to Chinese linguistic and societal backgrounds, as well as from existing NLP tasks in Chinese and English. To improve the evaluation robustness, we provide a diversified set of $5$ instructions with the same semantics for each task. Considering the potential data leakage issue of LLM benchmarks, we split two halves of $100$ input-output pairs in each task into private and public partitions, and only test and release the public split which contains one instruction variant. In sum, there are $45,000$ human-annotated [instruction, input, output] instances produced in CIF-Bench, as suggested in Table 1. In addition, we provide detailed instructions for all the tasks in § A.1.
@@
3.1 Data Collection
Annotation Protocol
CIF-Bench is designed for the extensive evaluation of Chinese comprehension and generation capabilities in LLMs, particularly focusing on aspects such as creative generation and linguistic abilities that existing benchmarks, such as C-Eval Huang et al. (2023b) and C-MMLU Li et al. (2023), struggle to assess. First, we select 113 diverse existing English NLP tasks, as shown in Table 1 from Super Natural Instructions (SNIWang et al. (2022b) and other research work (full list in § A.1). We then describe these task instructions in Chinese and a semantically balanced distributed subset from each original English NLP task as the Public split of CIF-Bench. We further ask expert native Chinese speakers, who minimally have undergraduate degrees, to annotate 100 samples per task based on the translated task instructions. These samples are further deduplicated according to their semantic embeddings. We finally select 50 samples per task as the Private split of CIF-Bench, to guarantee each sample’s validity and the balanced label distribution of each task.
To further improve CIF-Bench’s task diversity, we create 37 additional new tasks and state the related Chinese instructions. Specifically, we focus on adding Chinese tasks about Creative Natural Language Generation, Traditional Chinese, and Complex Role-Playing Text Games. We ask the expert native speakers to annotate 200 samples per task based on the translated task instructions. These samples are deduplicated and we further select the Public and Private split from it. Each task is further annotated with 4 Private paraphrased instructions to test whether LLMs understand the Chinese instructions’ meanings or overfit to the instructions in the Public split. Each sample and instruction is manually guaranteed or written by at least one of the authors to make sure that CIF-Bench is reliable.
@@ 3.2 Task Category
Whilst diverse tasks are provided in CIF-Bench, it would be difficult to analyze the extensive scores from all of the tasks. By reviewing and summarizing the existing NLP tasks and instruction-following benchmarks, we accordingly categorize the $150$ tasks into 20 basic types in a multi-label fashion (i.e., a task can be belong to more than one category). Each category consists of $2$ to $36$ tasks and the quantity distribution is revealed in Figure 2. Other than the $36$ “commonsense” tasks requiring a wide-ranging knowledge base, there are two dominant categories that aim to challenge the logical reasoning abilities of LLMs in CIF-Bench, including $30$ ”natural language inference (NLI)” and $29$ “reasoning” tasks. In particular, there are 18 tasks designed to require knowledge of unique Chinese cultural contexts. We describe the definition of each category and the task numbers in § A.2.
@@ 3.3 Task-based Automatic Evaluation
One core dilemma in evaluating the open-ended instruction-following capabilities of LLMs is that model predictions are hard to verify even with reference answers. For instance, it is intractable to handcraft regex rules to extract the predictions from LLMs for the extensive number of tasks, since the answers could be expressed in various formats, or drowned in redundant contexts like reasoning progress. Inspired by G-Eval Liu et al. (2023), we leverage OpenAI’s GPT-4222https://openai.com/gpt-4 as a relatively reliable evaluator for multi-class classification, multi-label classification, and creative generation tasks, to overcome such issues. The GPT-4 evaluator is prompted to assess the outputs according to the given task instruction and the input-output reference, as shown by the example in  Figure 3 and the full list of evaluation prompts in § A.1. For the answers that can be evaluated with semantic similarity, we use a lightweight multilingual encoder, BLEURT Sellam et al. (2020), to measure the relevance between the reference and LLM output.
@@ 4 Experiments
Model Name Score Differencenormal-$\uparrow$ Model Name Score Differencenormal-$\uparrow$
@@ 5 Results Analysis
In addition, we provide finer-grained score aggregation to further analyze the challenging task categories (n.b., most bilingual LLMs perform poorly on tasks in code, summarization, and translation categories). We point out that parallel data of Chinese and minor languages and Chinese-commented code corpora are still a scarce resource, which might lead to their poor performance on CIF-Bench’s code and translation categories. Additionally, we assume that Chinese and English bilingual LLMs, although a major branch of multilingual LLM, do not significantly benefit LLMs’ capacity to deal with minor-language-related tasks. Part of the tasks in CIF-Bench’s summarization category are very challenging, combining counterfactual reasoning and empathy estimation (i.e., task 126 and task 132 referring to Appendix A.1). Thereby, the bilingual LLMs’ poor performance on CIF-Bench’s summarization category is understandable. Detailed category-based scores on the public split are available in Table 13 in § B for further analysis.
@@
Model Name Var. Differencenormal-$\downarrow$ Model Name Var. Differencenormal-$\downarrow$
@@
To verify the annotation quality and reliability , we invite 3 annotators with expert-level NLP research backgrounds to assess the model outputs in public split with the same task-level instruction. The evaluation dimensions include: “Faithfulness”: human experts reflect on the absolute quality of a model’s output in a binary (yes/no) form. “Level of preference”: a 5-point Likert scale was provided to the experts to assess the relative quality of the model outputs. We randomly sample tasks according to the task category distribution, and pick three models performing differently in general, specifically Moss-Moon-003-sft ($0.399$), Baichuan-13B-Chat ($0.426$), and Qwen-72B-Chat ($0.589$). Considering the diverse and open-ended task, we first measure quality by comparing the pairwise agreement between each two annotators, reporting an average agreement of 0.49. Furthermore, we employ Cohen’s kappa (Ben-David, 2008) to measure inter-rater reliability, reporting an average of 0.3729 across the $153$ questions, implying that the results are substantially reliable. Specifically, the experts scored $0.4966$ on the dichotomous form yet $0.2492$ on the more varied options, suggesting that completing $153$ questions is challenging even for human experts. We further explore the correlation between the model prediction with human evaluation(Spearman’s $r=0.4043$), suggesting that most annotated were indeed truthful and the models can be relied upon to generate output for this task.
@@ A.1 Full List of Tasks and Evaluation
0 Negotiation Strategy Detection
SNI Wang et al. (2022b)
1 Grammar Error Correction
SNI Wang et al. (2022b)
2 Overlap Extraction
SNI Wang et al. (2022b)
3 Commonsense
SNI Wang et al. (2022b)
4 Data to Text
SNI Wang et al. (2022b)
5 Keyword Tagging
SNI Wang et al. (2022b)
6 Answerability Classification
SNI Wang et al. (2022b)
7 Dialogue Act Recognition
SNI Wang et al. (2022b)
8 Cause Effect Classification
SNI Wang et al. (2022b)
9 Question Rewriting
SNI Wang et al. (2022b)
10 Textual Entailment
SNI Wang et al. (2022b)
11 Coreference Resolution
SNI Wang et al. (2022b)
12 Title Generation
SNI Wang et al. (2022b)
13 Entity Relation Classification
SNI Wang et al. (2022b)
14 Punctuation Error Detection
SNI Wang et al. (2022b)
15 Style Transfer
SNI Wang et al. (2022b)
16 Sentence Expansion
SNI Wang et al. (2022b)
17 Poem Generation
SNI Wang et al. (2022b)
18 Discourse Relation Classification
SNI Wang et al. (2022b)
19 Mathematics
SNI Wang et al. (2022b)
20 Text Simplification
SNI Wang et al. (2022b)
21 Sentence Compression
SNI Wang et al. (2022b)
22 Spelling Error Detection
SNI Wang et al. (2022b)
23 Irony Detection
SNI Wang et al. (2022b)
24 Number Conversion
SNI Wang et al. (2022b)
25 Word Relation Classification
SNI Wang et al. (2022b)
26 Paraphrasing
SNI Wang et al. (2022b)
27 Grammar Error Detection
SNI Wang et al. (2022b)
28 Text Matching
SNI Wang et al. (2022b)
29 Fill in The Blank
SNI Wang et al. (2022b)
30 Speaker Relation Classification
SNI Wang et al. (2022b)
31 Entity Generation
SNI Wang et al. (2022b)
32 Summarization
SNI Wang et al. (2022b)
33 Spam Classification
SNI Wang et al. (2022b)
34 Stereotype Detection
SNI Wang et al. (2022b)
35 Dialogue State Tracking
SNI Wang et al. (2022b)
36 Dialogue State Tracking
SNI Wang et al. (2022b)
37 Sentence Perturbation
SNI Wang et al. (2022b)
38 Text Quality Evaluation
SNI Wang et al. (2022b)
39 Linguistic Probing
SNI Wang et al. (2022b)
40 Information Extraction
SNI Wang et al. (2022b)
41 Emotion Prediction
SNI Wang et al. (2022b)
42 Discourse Connective Identification
SNI Wang et al. (2022b)
43 Question Generation
SNI Wang et al. (2022b)
44 Stance Detection
SNI Wang et al. (2022b)
45 Sentiment Analysis
SNI Wang et al. (2022b)
46 Story Composition
SNI Wang et al. (2022b)
47 Program Execution
SNI Wang et al. (2022b)
48 Gender Classification
SNI Wang et al. (2022b)
49 Named Entity Recognition
SNI Wang et al. (2022b)
50 Toxic Language Detection
SNI Wang et al. (2022b)
51 Question Decomposition
SNI Wang et al. (2022b)
52 Sentence Ordering
SNI Wang et al. (2022b)
53 Text to Code
SNI Wang et al. (2022b)
54 Fact Verification
SNI Wang et al. (2022b)
55 Speaker Identification
SNI Wang et al. (2022b)
56 Answer Verification
SNI Wang et al. (2022b)
57 Wrong Candidate Generation
SNI Wang et al. (2022b)
58 Dialogue Generation
SNI Wang et al. (2022b)
59 Text Completion
SNI Wang et al. (2022b)
60 Pos Tagging
SNI Wang et al. (2022b)
@@ 60 Pos Tagging
61 Explanation
SNI Wang et al. (2022b)
62 Sentence Composition
SNI Wang et al. (2022b)
63 Question Understanding
SNI Wang et al. (2022b)
64 Intent Identification
SNI Wang et al. (2022b)
65 Word Semantics
SNI Wang et al. (2022b)
66 Code to Text
SNI Wang et al. (2022b)
67 Preposition Prediction
SNI Wang et al. (2022b)
68 Text Categorization
SNI Wang et al. (2022b)
69 Question Answering
SNI Wang et al. (2022b)
70 Commonsense Classification
N/A
71 Ancient Chinese Poem Retrieval
N/A
72 Ancient Chinese Translation
N/A
73 Chinese Rhyme Detection
N/A
74 Nationality Detection
N/A
75 Region Detection
N/A
76 Chinese Idiom Explanation
N/A
77 Name Allusion Detection
N/A
78 Chinese Ambiguity Sentence Location
N/A
79 Chinese Winograd Schema Challenge
FewCLUE Xu et al. (2021)
80 Chinese Modern Abbreviation Explanation
Wikipedia Wikipedia (2024)
81 Chinese Epigraph Detection
N/A
82 Chinese Dialect Translation
N/A
83 Chinese Attractions List
N/A
85 Chinese Typo Categorization
N/A
86 Chinese Fiction Characteristic Detection
N/A
87 Chinese Figurative Detection
N/A
88 Chinese Metaphor Explanation
N/A
89 Chinese Medicine Detection
N/A
90 Chinese Pinyin Detection
N/A
91 Chinese Wubi Written
N/A
92 Intimacy Score Prediction
Pei and Jurgens (2020)
93 Sentence Level Uncertainty Judgement
Pei and Jurgens (2021)
94 Chinese Relative Identification
N/A
96 Chinese Heteronomous Language Detection
N/A
97 Code Debug
https://blog.csdn.net
98 Code Translate
https://leetcode.cn
99 Function Explanation
https://www.liaoxuefeng.com/
100 Bias Detoxication
CORGI-PM Zhang et al. (2023a)
101 MultiLabel Chinese Humor Categorization
Tseng et al. (2020)
102 Legal Term Retrieval
N/A
103 Patronizing Condescending Multilabel
Perez-Almendros et al. (2022)
104 CommonSense Explanation
Wang et al. (2020)
105 Event Type Detection
MUSIED (Xi et al., 2022)
106 Argument Mining
N/A
107 Theory of Mind
Big-Bench Theory of Mind (bench authors, 2023)
108 Game Playing
Big-Bench Language Games (bench authors, 2023)
109 IQ Test
Huang et al. (2023a)
110 Joke Explanation
N/A
111 Role Playing
TRPG https://bilibili.com Côté et al. (2018)
112 Text De-Identification
N/A
113 Outline Generation
N/A
114 Pros Cons Listing
N/A
115 Joke Telling
N/A
116 Affordance
N/A

### v2

@@ Abstract
The advancement of large language models (LLMs) has enhanced the ability to generalize across a wide range of unseen natural language processing (NLP) tasks through instruction-following. Yet, their effectiveness often diminishes in less-trained languages like Chinese, exacerbated by biased evaluations from data leakage, casting doubt on their true generalizability to new linguistic territories. In response, we introduce the Chinese Instruction-Following Benchmark (CIF-Bench), designed to evaluate the zero-shot generalizability of LLMs to the Chinese language. CIF-Bench comprises 150 tasks and 15,000 input-output pairs, developed by native speakers to test complex reasoning and Chinese cultural nuances across 20 categories. To mitigate data contamination, we release only half of the dataset publicly, with the remainder kept private, and introduce diversified instructions to minimize score variance, totaling 45,000 data instances. Our evaluation of $28$ selected LLMs reveals a noticeable performance gap, with the best model scoring only 52.9%, highlighting the limitations of LLMs in less familiar language and task contexts. This work aims to uncover the current limitations of LLMs in handling Chinese tasks, pushing towards the development of more culturally informed and linguistically diverse models with the released data and benchmark [^1] https://yizhilll.github.io/CIF-Bench/.
@@ 1 Introduction
The landscape of natural language processing (NLP) has been dramatically reshaped by the emergence of large language models (LLMs), which have demonstrated an ability to generalize across unseen NLP tasks, often showcased through the framework of instruction-following Mishra et al. (2021); Sanh et al. (2021); Wei et al. (2021). Despite these advances, skepticism remains regarding the transferability of this instruction-following capability, particularly in multilingual contexts. The models perform worse when switching to Chinese due to the prevalence of English training data Huang et al. (2023b); Zhang et al. (2023b), as figured in Fig. 1. This concern is exacerbated by observations that benchmarks designed to assess the capabilities of LLMs may inadvertently suffer from biased evaluations due to data leakage Sainz et al. (2023), particularly when web-scale datasets are employed to enhance model generalizability Raffel et al. (2023). Such observations raise a critical question: While the generalizability of LLMs appears intriguing, do these models face significant challenges when evaluated on private and diversified instruction-formatted tasks in less common language contexts?
@@ 2 Related Work
Large-scale pre-trained language models have been found that they can generalize across unseen tasks by fine-tuned on formatted task instructions (Khashabi et al., 2020; Mishra et al., 2021; Wei et al., 2021; Sanh et al., 2021). Early studies attempt to fine-tune and evaluate such a capability in a few-shot manner by providing input-output examples Ye et al. (2021); Mishra et al. (2021). Following that, another line of research Bach et al. (2022); Wang et al. (2022b)  Bai et al. (2024) improves the evaluation reliability from the perspective of scaling the task quantity and providing well-defined corresponding instructions. A more recent concurrent work FollowBench proposes to craft multiple instructions for a single task to evaluate the LLMs, similar to CIF-Bench. A core distinction between CIF-Bench and the FollowBench is that we focus on assessing whether models can stably perform given diversely expressed, but semantically identical instructions, while FollowBench aims to extend the basic instruction with different additional requirements.
@@ 3 The Challenging Chinese Instruction-Following Benchmark
We define a total of 150 curated tasks, constructed according to Chinese linguistic and societal backgrounds, as well as from existing NLP tasks in Chinese and English. To improve the evaluation robustness, we provide a diversified set of $5$ instructions with the same semantics for each task. Considering the potential data leakage issue of LLM benchmarks, we split two halves of $100$ input-output pairs in each task into private and public partitions, and only test and release the public split which contains one instruction variant. In sum, there are $45,000$ human-annotated [instruction, input, output] instances produced in CIF-Bench, as suggested in Table 1. In addition, we provide detailed instructions for all the tasks in Appendix A.1.
@@
3.1 Data Collection.
Collecting Sources.
CIF-Bench is designed for the extensive evaluation of Chinese comprehension and generation capabilities in LLMs, particularly focusing on aspects such as creative generation and linguistic abilities that existing benchmarks, such as C-Eval Huang et al. (2023b) and C-MMLU Li et al. (2023), struggle to assess. First, we select 113 diverse existing English NLP tasks, as shown in Table 1 from Super Natural Instructions (SNIWang et al. (2022b) and other research work (full list in Appendix A.1). We then describe these task instructions in Chinese and a semantically balanced distributed subset from each original English NLP task as the Public split of CIF-Bench. We further ask expert native Chinese speakers, who minimally have undergraduate degrees, to annotate 100 samples per task based on the translated task instructions. These samples are further deduplicated according to their semantic embeddings. We finally select 50 samples per task as the Private split of CIF-Bench, to guarantee each sample’s validity and the balanced label distribution of each task.
Annotation Protocol.
To be specific, we set up a robust three-stage pipeline in our annotation process. In stage 1, to ensure high annotation quality, we hire native speakers with college backgrounds to annotate the data samples in the form of triplet <instruction, input, output> in cooperated with the annotation platform Stardust [^2] https://stardust.ai. In stage 2, the data annotation specialists from the platform conduct a second round of checking on the quality of the samples. The specialists first use the GPT-4 as an auxiliary verification, and the samples scored lower than 6 out of 10 would be directly deleted. The specialists then manually check on the rest of the samples and deleted the unqualified ones. Next, annotators from the stage 1 would continue the annotation until collecting 100 input-output pairs per task. The specialists also check on the distribution of the labels and answers, to avoid similar input-output pairs for the task. In stage 3, four researchers with NLP backgrounds conduct a final check by inspecting randomly sampled 20 data points from the 150 tasks. If one of the samples does not satisfy the annotation requirements, the task will be returned to the beginning of the annotation pipeline until it passes verification. Such a pipeline of three stages costs approximately \$24K.
Detailed Categories.
To further improve CIF-Bench’s task diversity, we create 37 additional new tasks and state the related Chinese instructions. Specifically, we focus on adding Chinese tasks about Creative Natural Language Generation, Traditional Chinese, and Complex Role-Playing Text Games. We ask the expert native speakers to annotate 200 samples per task based on the translated task instructions. These samples are deduplicated and we further select the Public and Private split from it. Each task is further annotated with 4 Private paraphrased instructions to test whether LLMs understand the Chinese instructions’ meanings or overfit to the instructions in the Public split. Each sample and instruction is manually verified or written by the authors to make sure that CIF-Bench is reliable.
@@ 3.2 Task Category
Whilst diverse tasks are provided in CIF-Bench, it would be difficult to analyze the extensive scores from all of the tasks. By reviewing and summarizing the existing NLP tasks and instruction-following benchmarks, we accordingly categorize the $150$ tasks into 20 basic types in a multi-label fashion (i.e., a task can be belong to more than one category). Each category consists of $2$ to $36$ tasks and the quantity distribution is revealed in Figure 2. Other than the $36$ “commonsense” tasks requiring a wide-ranging knowledge base, there are two dominant categories that aim to challenge the logical reasoning abilities of LLMs in CIF-Bench, including $30$ ”natural language inference (NLI)” and $29$ “reasoning” tasks. In particular, there are 18 tasks designed to require knowledge of unique Chinese cultural contexts. We describe the definition of each category and the task numbers in Appendix A.2.
@@ 3.3 Task-based Automatic Evaluation
One core dilemma in evaluating the open-ended instruction-following capabilities of LLMs is that model predictions are hard to verify even with reference answers. For instance, it is intractable to handcraft regex rules to extract the predictions from LLMs for the extensive number of tasks, since the answers could be expressed in various formats, or drowned in redundant contexts like reasoning progress. Inspired by G-Eval Liu et al. (2023), we leverage OpenAI’s GPT-4333https://openai.com/gpt-4 as a relatively reliable evaluator for multi-class classification, multi-label classification, and creative generation tasks, to overcome such issues. The GPT-4 evaluator is prompted to assess the outputs according to the given task instruction and the input-output reference, as shown by the example in  Figure 3 and the full list of evaluation prompts in Appendix A.1. the remaining tasks that can be evaluated by the semantic similarity between the golden reference and model output, we use a lightweight multilingual encoder, BLEURT Sellam et al. (2020), to measure the relevance between the reference and LLM output.
@@ 4 Experiments
[…]
@@ 5 Results Analysis
In addition, we provide finer-grained score aggregation to further analyze the challenging task categories (n.b., most bilingual LLMs perform poorly on tasks in code, summarization, and translation categories). In the code category, the models might misunderstand the semantics expressed in Chinese for the newly defined variable or function. Specifically, models usually perform poorly in a new “programming language” environment that requires the model to understand restricted actions. As for summarization tasks, models could misinterpret the instruction, eg. models sometimes consider the instruction “modify the input into a more friendly expression to non-native speakers” as a Chinese-English translation task and might provide redundant explanations even if not required by the instructions and hence will cause large semantic distances to the golden reference. We point out that Chinese-commented code corpora and parallel translation data of Chinese and other languages are still scarce resources, which might lead to their poor performance on CIF-Bench’s code and translation categories. Additionally, we assume that Chinese and English bilingual LLMs, although a major branch of multilingual LLM, do not significantly benefit LLMs’ capacity to deal with minor-language-related tasks. Part of the tasks in CIF-Bench’s summarization category are very challenging, combining counterfactual reasoning and empathy estimation (i.e., task 125 and task 131 referring to Appendix A.1). Thereby, the bilingual LLMs’ poor performance on CIF-Bench’s summarization category is understandable. Detailed category-based scores on the public split are available in Table 13 in Appendix B for further analysis.
@@
[…]
@@
To verify the annotation quality and reliability , we invite 3 annotators with expert-level NLP research backgrounds to assess the model outputs in public split with the same task-level instruction. The evaluation dimensions include: “Faithfulness”: human experts reflect on the absolute quality of a model’s output in a binary (yes/no) form. “Level of preference”: a 5-point Likert scale was provided to the experts to assess the relative quality of the model outputs. We randomly sample tasks according to the task category distribution, and pick three models performing differently in general, specifically Moss-Moon-003-sft ($0.399$), Baichuan-13B-Chat ($0.426$), and Qwen-72B-Chat ($0.589$). Considering the diverse and open-ended task, we first measure quality by comparing the pairwise agreement between two annotators, reporting an average agreement of 0.49. Furthermore, we employ Cohen’s kappa (Ben-David, 2008) to measure inter-rater reliability, reporting an average of 0.3729 across the $153$ questions, implying that the results are substantially reliable. Specifically, the experts scored $0.4966$ on the dichotomous form yet $0.2492$ on the more varied options, suggesting that completing $153$ questions is challenging even for human experts. We further explore the correlation between the model prediction with human evaluation(Spearman’s $r=0.4043$), suggesting that most annotated were indeed truthful and the models can be relied upon to generate output for this task.
@@ 6 Conclusion
Acknowledgements
Yizhi Li and Xingwei Qu are Ph.D. students funded by the Department of Computer Science, University of Manchester, UK.
@@ A.1 Full List of Tasks and Evaluation
0 Negotiation Strategy Detection SNI Wang et al. (2022b)
1 Grammar Error Correction SNI Wang et al. (2022b)
2 Overlap Extraction SNI Wang et al. (2022b)
3 Commonsense SNI Wang et al. (2022b)
4 Data to Text SNI Wang et al. (2022b)
5 Keyword Tagging SNI Wang et al. (2022b)
6 Answerability Classification SNI Wang et al. (2022b)
7 Dialogue Act Recognition SNI Wang et al. (2022b)
8 Cause Effect Classification SNI Wang et al. (2022b)
9 Question Rewriting SNI Wang et al. (2022b)
10 Textual Entailment SNI Wang et al. (2022b)
11 Coreference Resolution SNI Wang et al. (2022b)
12 Title Generation SNI Wang et al. (2022b)
13 Entity Relation Classification SNI Wang et al. (2022b)
14 Punctuation Error Detection SNI Wang et al. (2022b)
15 Style Transfer SNI Wang et al. (2022b)
16 Sentence Expansion SNI Wang et al. (2022b)
17 Poem Generation SNI Wang et al. (2022b)
18 Discourse Relation Classification SNI Wang et al. (2022b)
19 Mathematics SNI Wang et al. (2022b)
20 Text Simplification SNI Wang et al. (2022b)
21 Sentence Compression SNI Wang et al. (2022b)
22 Spelling Error Detection SNI Wang et al. (2022b)
23 Irony Detection SNI Wang et al. (2022b)
24 Number Conversion SNI Wang et al. (2022b)
25 Word Relation Classification SNI Wang et al. (2022b)
26 Paraphrasing SNI Wang et al. (2022b)
27 Grammar Error Detection SNI Wang et al. (2022b)
28 Text Matching SNI Wang et al. (2022b)
29 Fill in The Blank SNI Wang et al. (2022b)
30 Speaker Relation Classification SNI Wang et al. (2022b)
31 Entity Generation SNI Wang et al. (2022b)
32 Summarization SNI Wang et al. (2022b)
33 Spam Classification SNI Wang et al. (2022b)
34 Stereotype Detection SNI Wang et al. (2022b)
35 Dialogue State Tracking SNI Wang et al. (2022b)
36 Dialogue State Tracking SNI Wang et al. (2022b)
37 Sentence Perturbation SNI Wang et al. (2022b)
38 Text Quality Evaluation SNI Wang et al. (2022b)
39 Linguistic Probing SNI Wang et al. (2022b)
40 Information Extraction SNI Wang et al. (2022b)
41 Emotion Prediction SNI Wang et al. (2022b)
42 Discourse Connective Identification SNI Wang et al. (2022b)
43 Question Generation SNI Wang et al. (2022b)
44 Stance Detection SNI Wang et al. (2022b)
45 Sentiment Analysis SNI Wang et al. (2022b)
46 Story Composition SNI Wang et al. (2022b)
47 Program Execution SNI Wang et al. (2022b)
48 Gender Classification SNI Wang et al. (2022b)
49 Named Entity Recognition SNI Wang et al. (2022b)
50 Toxic Language Detection SNI Wang et al. (2022b)
51 Question Decomposition SNI Wang et al. (2022b)
52 Sentence Ordering SNI Wang et al. (2022b)
53 Text to Code SNI Wang et al. (2022b)
54 Fact Verification SNI Wang et al. (2022b)
55 Speaker Identification SNI Wang et al. (2022b)
56 Answer Verification SNI Wang et al. (2022b)
57 Wrong Candidate Generation SNI Wang et al. (2022b)
58 Dialogue Generation SNI Wang et al. (2022b)
59 Text Completion SNI Wang et al. (2022b)
60 Pos Tagging SNI Wang et al. (2022b)
@@
61 Explanation SNI Wang et al. (2022b)
62 Sentence Composition SNI Wang et al. (2022b)
63 Question Understanding SNI Wang et al. (2022b)
64 Intent Identification SNI Wang et al. (2022b)
65 Word Semantics SNI Wang et al. (2022b)
66 Code to Text SNI Wang et al. (2022b)
67 Preposition Prediction SNI Wang et al. (2022b)
68 Text Categorization SNI Wang et al. (2022b)
69 Question Answering SNI Wang et al. (2022b)
70 Commonsense Classification N/A
71 Ancient Chinese Poem Retrieval N/A
72 Ancient Chinese Translation N/A
73 Chinese Rhyme Detection N/A
74 Nationality Detection N/A
75 Region Detection N/A
76 Chinese Idiom Explanation N/A
77 Name Allusion Detection N/A
78 Chinese Ambiguity Sentence Location N/A
79 Chinese Winograd Schema Challenge FewCLUE Xu et al. (2021)
80 Chinese Modern Abbreviation Explanation Wikipedia Wikipedia (2024)
81 Chinese Epigraph Detection N/A
82 Chinese Dialect Translation N/A
83 Chinese Attractions List N/A
85 Chinese Typo Categorization N/A
86 Chinese Fiction Characteristic Detection N/A
87 Chinese Figurative Detection N/A
88 Chinese Metaphor Explanation N/A
89 Chinese Medicine Detection N/A
90 Chinese Pinyin Detection N/A
91 Chinese Wubi Written N/A
92 Intimacy Score Prediction Pei and Jurgens (2020)
93 Sentence Level Uncertainty Judgement Pei and Jurgens (2021)
94 Chinese Relative Identification N/A
96 Chinese Heteronomous Language Detection N/A
97 Code Debug https://blog.csdn.net
98 Code Translate https://leetcode.cn
99 Function Explanation https://www.liaoxuefeng.com/
100 Bias Detoxication CORGI-PM Zhang et al. (2023a)
101 MultiLabel Chinese Humor Categorization Tseng et al. (2020)
102 Legal Term Retrieval N/A
103 Patronizing Condescending Multilabel Perez-Almendros et al. (2022)
104 CommonSense Explanation Wang et al. (2020)
105 Event Type Detection MUSIED (Xi et al., 2022)
106 Argument Mining N/A
107 Theory of Mind Big-Bench Theory of Mind (bench authors, 2023)
108 Game Playing Big-Bench Language Games (bench authors, 2023)
109 IQ Test Huang et al. (2023a)
110 Joke Explanation N/A
111 Role Playing TRPG https://bilibili.com Côté et al. (2018)
112 Text De-Identification N/A
113 Outline Generation N/A
114 Pros Cons Listing N/A
115 Joke Telling N/A
116 Affordance N/A